# Zero-Shot Trajectory Planning for Signal Temporal Logic Tasks

### v1

@@ D.3 Experimental Parameter Settings
  • Sampling Factor ($\eta$): Used when computing the robustness value of trajectories, as described in Section D.2.
Env $N_{\mathrm{max}}$ $H$ $N$ $\gamma$ $\eta$
Maze2D-Umaze 1 128 64 0.8 8
Maze2D-Medium 1 256 256 0.9 8
Maze2D-Large 1 384 256 1.1 8
Custom-Built Env 1 16&32 64 1 1
Table D.1: Parameters Used in the Experiments
Appendix E More Cases
We visualized some of the experimental results. The actual execution trajectories for some successful cases (where the STL tasks were satisfied by the execution trajectories) are shown in Figure E.1 to Figure E.3. For some failed cases (where the STL tasks were not satisfied), the trajectories planned by our algorithm are shown in Figure E.4.
Notably, by analyzing the failed cases, we identified that the primary reason for execution failure is that the trajectories generated by the diffusion model significantly violated system dynamics, such as colliding with obstacles in the environment or having excessively large distances between consecutive states. To further enhance the actual execution success rate, our method can be integrated with some receding horizon control methods Zhou et al. [2024a] or online replanning strategies Zhou et al. [2024b]. This extension will be explored in our future work.
Figure E.1: Successful Execution Trajectories in Maze2D-large Environment under STL Template 3: $\mathrm{F}_{\mathrm{I}_{1}}\mu_{1}\wedge(\neg\mu_{1}\mathrm{U}_{\mathrm{I}_{1}}\mu_{2})$. The task requires the agent to eventually reach region $\mu_{1}$ within a given time interval, but before reaching $\mu_{1}$, it must first visit region $\mu_{2}$.
Figure E.2: Successful Execution Trajectories in Maze2D-medium Environment under STL Template 4: $\mathrm{F}_{\mathrm{I}_{1}}(\mu_{1}\wedge(\mathrm{F}_{\mathrm{I}_{2}}(\mu_{2}\wedge\mathrm{F}_{\mathrm{I}_{3}}(\mu_{3}\wedge\mathrm{F}_{\mathrm{I}_{4}}(\mu_{4})))))$. The task requires the agent to sequentially visit regions $\mu_{1},\mu_{2},\mu_{3},\mu_{4}$ within the given time intervals.
Figure E.3: Successful Execution Trajectories in Maze2D-umaze Environment under STL Template 5: $\mathrm{F}_{\mathrm{I}_{1}}(\mu_{1}\wedge(\mathrm{F}_{\mathrm{I}_{2}}(\mu_{2}\wedge\mathrm{F}_{\mathrm{I}_{3}}(\mu_{3}))))\wedge\mathrm{G}(\neg\mu_{4})$. The task requires the agent to sequentially visit regions $\mu_{1},\mu_{2},\mu_{3}$ within the given time intervals and always avoid the region $\mu_{4}$.
Figure E.4: Planned Trajectories in Some Failure Cases.

### v2

@@ F.7.2 Experimental Parameter Settings
  • Sampling Factor ($\eta$): Used when computing the robustness value of trajectories, as described in Section F.7.1.
  • Control Frequency ($k$): The number of low-level control updates executed per planning step during runtime, as described in Section 3.5.
Appendix G Proofs
G.1 Proof of Lemma 1
Proof.
We prove the statement by structural induction on $\varphi$. Throughout, we assume discrete time $t\in\mathbb{Z}_{\geq 0}$ and use the shorthand $\mathbf{s}_{0}\vDash\mathcal{R}(a,b,\mu)\Leftrightarrow\exists t\in[a,b]:\mathbf{x}_{t}\vDash\mu$ and $\mathbf{s}_{0}\vDash\mathcal{I}(a,b,\mu)\Leftrightarrow\forall t\in[a,b]:\mathbf{x}_{t}\vDash\mu$ (cf. (3)-(4)).
Base cases.
(i) $\varphi=\mathrm{F}_{[a,b]}\mu$. The decomposition yields $\mathbb{P}_{\varphi}=\{\mathcal{R}(\lambda,\lambda,\mu)\}$, $\mathbb{T}_{\varphi}=\{\lambda\in[a,b]\}$, $\Lambda_{\varphi}=\{\lambda\}$. By the Boolean semantics of $\mathrm{F}$ (Eq. (C.4)),
$\mathbf{s}_{0}\vDash\mathrm{F}_{[a,b]}\mu\ \Leftrightarrow\ \exists\lambda\in[a,b]\text{ s.t. }\mathbf{x}_{\lambda}\vDash\mu\ \Leftrightarrow\ \exists\lambda\in[a,b]\text{ s.t. }\mathbf{s}_{0}\vDash\mathcal{R}(\lambda,\lambda,\mu),$
which is exactly the desired form with $\boldsymbol{\lambda}=[\lambda]\in\mathcal{F}_{\varphi}$.
(ii) $\varphi=\mathrm{G}_{[a,b]}\mu$. The decomposition yields $\mathbb{P}_{\varphi}=\{\mathcal{I}(a,b,\mu)\}$ and $\Lambda_{\varphi}=\mathbb{T}_{\varphi}=\varnothing$. By Eq. (C.5),
$\mathbf{s}_{0}\vDash\mathrm{G}_{[a,b]}\mu\ \Leftrightarrow\ \forall t\in[a,b]:\mathbf{x}_{t}\vDash\mu\ \Leftrightarrow\ \mathbf{s}_{0}\vDash\mathcal{I}(a,b,\mu),$
which matches the target equivalence (no time variables).
(iii) $\varphi=\mu_{1}\mathrm{U}_{[a,b]}\mu_{2}$. The decomposition yields $\mathbb{P}_{\varphi}=\{\mathcal{I}(0,\lambda,\mu_{1}),\mathcal{R}(\lambda,\lambda,\mu_{2})\}$, $\mathbb{T}_{\varphi}=\{\lambda\in[a,b]\}$, $\Lambda_{\varphi}=\{\lambda\}$. By Eq. (C.6),
$\mathbf{s}_{0}\vDash\mu_{1}\mathrm{U}_{[a,b]}\mu_{2}$ $\Leftrightarrow\exists\lambda\in[a,b]\!:\ \mathbf{x}_{\lambda}\vDash\mu_{2}\ \wedge\ \forall t\in[0,\lambda]:\mathbf{x}_{t}\vDash\mu_{1}$
$\Leftrightarrow\exists\lambda\in[a,b]\!:\ \mathbf{s}_{0}\vDash\mathcal{R}(\lambda,\lambda,\mu_{2})\ \wedge\ \mathbf{s}_{0}\vDash\mathcal{I}(0,\lambda,\mu_{1}),$
again yielding the target form with $\boldsymbol{\lambda}=[\lambda]\in\mathcal{F}_{\varphi}$.
Inductive steps. For each constructor below we assume the lemma holds for the immediate subformula(e).
(1) Conjunction $\varphi=\varphi_{1}\wedge\varphi_{2}$. By semantics, $\mathbf{s}_{0}\vDash\varphi\Leftrightarrow(\mathbf{s}_{0}\vDash\varphi_{1}\ \wedge\ \mathbf{s}_{0}\vDash\varphi_{2})$. The decomposition uses disjoint unions $\mathbb{P}_{\varphi}=\mathbb{P}_{\varphi_{1}}\uplus\mathbb{P}_{\varphi_{2}},\ \mathbb{T}_{\varphi}=\mathbb{T}_{\varphi_{1}}\uplus\mathbb{T}_{\varphi_{2}},\ \Lambda_{\varphi}=\Lambda_{\varphi_{1}}\uplus\Lambda_{\varphi_{2}},$ implicitly $\alpha$-renaming to avoid clashes. Since each constraint in $\mathbb{T}_{\varphi_{i}}$ involves a single variable, the feasible set factors: $\mathcal{F}_{\varphi}=\mathcal{F}_{\varphi_{1}}\times\mathcal{F}_{\varphi_{2}}.$ By the IH on $\varphi_{1},\varphi_{2}$ and the fact that every progress in $\mathbb{P}_{\varphi}$ depends only on its own component of $\boldsymbol{\lambda}=[\boldsymbol{\lambda}_{1},\boldsymbol{\lambda}_{2}]$, we obtain
$\mathbf{s}_{0}\vDash\varphi\ \Leftrightarrow\ \exists\boldsymbol{\lambda}\in\mathcal{F}_{\varphi}\text{ s.t. }\forall\mathcal{P}\in\mathbb{P}_{\varphi}:\ \mathbf{s}_{0}\vDash\mathcal{P}(\boldsymbol{\lambda}).$
(2) Outer finally $\varphi^{\prime}=\mathrm{F}_{[a,b]}\varphi$. By Eq. (C.4), $\mathbf{s}_{0}\vDash\varphi^{\prime}\Leftrightarrow\exists\lambda\in[a,b]:\mathbf{s}_{\lambda}\vDash\varphi.$ Time-shift identity. For any progress $\mathcal{P}(c,d,\mu)$ and any $k\in\mathbb{Z}_{\geq 0}$,
$\mathbf{s}_{k}\vDash\mathcal{P}(c,d,\mu)\ \Leftrightarrow\ \mathbf{s}_{0}\vDash\mathcal{P}(c+k,d+k,\mu).$ (G.1)
This is immediate from the definitions of $\mathcal{R}$ and $\mathcal{I}$. The decomposition introduces a fresh $\lambda\in[a,b]$ and shifts every progress of $\varphi$ by $\lambda$; hence $\Lambda_{\varphi^{\prime}}=\Lambda_{\varphi}\uplus\{\lambda\}$ and $\mathcal{F}_{\varphi^{\prime}}=\{[\boldsymbol{\lambda},\lambda]\mid\boldsymbol{\lambda}\in\mathcal{F}_{\varphi},\ \lambda\in[a,b]\}.$ Applying the IH to $\varphi$ on the shifted signal $\mathbf{s}_{\lambda}$ and then using (G.1) yields
$\mathbf{s}_{0}\vDash\varphi^{\prime}\ \Leftrightarrow\ \exists\boldsymbol{\lambda}^{\prime}\in\mathcal{F}_{\varphi^{\prime}}\text{ s.t. }\forall\mathcal{P}^{\prime}\in\mathbb{P}_{\varphi^{\prime}}:\ \mathbf{s}_{0}\vDash\mathcal{P}^{\prime}(\boldsymbol{\lambda}^{\prime}).$
(The special case $\mathrm{F}_{[a,a]}$ is the same with a constant shift +a+\!a and no new variable.)
(3) Outer always $\varphi^{\prime}=\mathrm{G}_{[a,b]}\varphi$. In discrete time,
$\mathrm{G}_{[a,b]}\varphi\ \equiv\ \bigwedge_{k=a}^{b}\mathrm{F}_{[k,k]}\varphi,$ (G.2)
since $\mathbf{s}_{0}\vDash\mathrm{G}_{[a,b]}\varphi$ iff $\mathbf{s}_{k}\vDash\varphi$ for every integer $k\in[a,b]$. The decomposition makes an independent copy $(\mathbb{P}_{\varphi}^{(k)},\mathbb{T}_{\varphi}^{(k)},\Lambda_{\varphi}^{(k)})$ for each $k$ and shifts every progress by +k+k, then merges all copies:
$\mathbb{P}_{\varphi^{\prime}}=\bigcup_{k=a}^{b}\{\mathcal{P}(c_{\Lambda}+k,d_{\Lambda}+k,\mu)\mid\mathcal{P}(c_{\Lambda},d_{\Lambda},\mu)\in\mathbb{P}_{\varphi}^{(k)}\},\quad\Lambda_{\varphi^{\prime}}=\biguplus_{k=a}^{b}\Lambda_{\varphi}^{(k)},\quad\mathcal{F}_{\varphi^{\prime}}=\prod_{k=a}^{b}\mathcal{F}_{\varphi}.$
Applying item (2) to each $\mathrm{F}_{[k,k]}$ factor and using (G.1) yields the target form. Merging constant invariances. If $\mathcal{I}(c,d,\mu)\in\mathbb{P}_{\varphi}$ has constant endpoints, then its shifted copies $\{\mathcal{I}(c+k,d+k,\mu)\}_{k=a}^{b}$ are equivalent to the single $\mathcal{I}(c+a,d+b,\mu)$ because $\bigcup_{k=a}^{b}[c+k,d+k]=[c+a,d+b]$ in discrete time; hence merging preserves truth.
(4) Outer until $\varphi^{\prime}=\phi\mathrm{U}_{[a,b]}\varphi$. By assumption, $\phi$ contains no $\mathrm{F}$ or $\mathrm{U}$, thus its decomposition consists solely of invariance progresses with constant endpoints and $\Lambda_{\phi}=\mathbb{T}_{\phi}=\varnothing$. The decomposition of $\varphi^{\prime}$: introduce $\lambda\in[a,b]$, shift each progress of $\varphi$ by $\lambda$, and prolong every $\mathcal{I}(c,d,\mu)\in\mathbb{P}_{\phi}$ to $\mathcal{I}(c,d+\lambda,\mu)$. Formally,
$\mathbb{P}_{\varphi^{\prime}}=\{\mathcal{P}(c_{\Lambda}+\lambda,d_{\Lambda}+\lambda,\mu)\mid\mathcal{P}\in\mathbb{P}_{\varphi}\}\ \cup\ \{\mathcal{I}(c,d+\lambda,\mu)\mid\mathcal{I}(c,d,\mu)\in\mathbb{P}_{\phi}\},$
$\mathcal{F}_{\varphi^{\prime}}=\{[\boldsymbol{\lambda},\lambda]\mid\boldsymbol{\lambda}\in\mathcal{F}_{\varphi},\ \lambda\in[a,b]\}.$
By Eq. (C.6),
$\mathbf{s}_{0}\vDash\phi\mathrm{U}_{[a,b]}\varphi\ \Leftrightarrow\ \exists\lambda\in[a,b]\ \text{s.t.}\ \Big(\mathbf{s}_{\lambda}\vDash\varphi\ \wedge\ \forall\tau\in[0,\lambda]:\ \mathbf{s}_{\tau}\vDash\phi\Big).$ (G.3)
The first conjunct is handled by item (2) with the shift identity (G.1). For the second conjunct, note the following sliding-invariance identity: for any constant $c\leq d$,
$\big(\forall\tau\in[0,\lambda]:\ \mathbf{s}_{\tau}\vDash\mathcal{I}(c,d,\mu)\big)\ \Leftrightarrow\ \mathbf{s}_{0}\vDash\mathcal{I}(c,d+\lambda,\mu),$ (G.4)
because $\mathbf{s}_{\tau}\vDash\mathcal{I}(c,d,\mu)$ means $\forall t\in[c+\tau,d+\tau]:\mathbf{x}_{t}\vDash\mu$, and $\bigcup_{\tau=0}^{\lambda}[c+\tau,d+\tau]=[c,d+\lambda]$ in discrete time. Since $\mathbb{P}_{\phi}$ is a conjunction of such invariances, applying (G.4) to each member yields the prolonged invariances in $\mathbb{P}_{\varphi^{\prime}}$. Combining these observations with (G.3) gives
$\mathbf{s}_{0}\vDash\varphi^{\prime}\ \Leftrightarrow\ \exists\boldsymbol{\lambda}^{\prime}\in\mathcal{F}_{\varphi^{\prime}}\text{ s.t. }\forall\mathcal{P}^{\prime}\in\mathbb{P}_{\varphi^{\prime}}:\ \mathbf{s}_{0}\vDash\mathcal{P}^{\prime}(\boldsymbol{\lambda}^{\prime}).$
Conclusion. All constructors preserve the claimed equivalence. By structural induction, for any PNF STL formula without disjunctions,
$\mathbf{s}_{0}\vDash\varphi\ \Longleftrightarrow\ \exists\,\boldsymbol{\lambda}\in\mathcal{F}_{\varphi}\ \text{s.t.}\ \forall\mathcal{P}\in\mathbb{P}_{\varphi}:\ \mathbf{s}_{0}\vDash\mathcal{P}(\boldsymbol{\lambda}).$
If $\Lambda_{\varphi}=\varnothing$, then $\mathcal{F}_{\varphi}=\{\emptyset\}$ and the statement reduces to $\mathbf{s}_{0}\vDash\varphi\Longleftrightarrow\forall\mathcal{P}\in\mathbb{P}_{\varphi}:\ \mathbf{s}_{0}\vDash\mathcal{P}.$
G.2 Proof of Lemma 2
Proof.
[…]
Inductive invariants.
After initialization and after every successful extension by Algorithm 2 and UpdateConstraint, the following hold:
  • (I1)
    Feasibility is preserved: the current constraint set $\mathbb{T}$ admits at least one assignment of time variables (i.e., the feasible set is nonempty).
  • (I2)
    Committed reachabilities are honored: for every reachability progress $\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)$ that has been assigned a waypoint $(\mathbf{x}^{\prime},t^{\prime})$ and removed from $\mathbb{P}^{\mathcal{R}}_{\mathrm{rem}}$, $\mathbb{T}$ contains the inequalities $a_{\Lambda}\leq t^{\prime}$ and $b_{\Lambda}\geq t^{\prime}$. Hence any $\boldsymbol{\lambda}$ feasible for $\mathbb{T}$ satisfies $t^{\prime}\in[a_{\Lambda}(\boldsymbol{\lambda}),b_{\Lambda}(\boldsymbol{\lambda})]$.
  • (I3)
    Waypoint-level invariance consistency: let $\mathcal{I}(c,d_{\Lambda},\mu^{\prime})$ be any invariance progress whose start time has been determined (by preprocessing every $\mathcal{I}$ has a preceding $\mathcal{R}(c,c,\mu^{\prime})$; once that $\mathcal{R}$ is assigned at time $c=t^{*}$, the start is fixed). Whenever a new waypoint $(\mathbf{x}^{\prime},t^{\prime})$ is appended with $\mathbf{x}^{\prime}\nvDash\mu^{\prime}$, Algorithm 2 computes a conflict interval $\mathcal{O}$ and UpdateConstraint adds $d_{\Lambda}<t^{\prime}$; consequently, no feasible $\boldsymbol{\lambda}$ can make the interval $c,d_{\Lambda}(\boldsymbol{\lambda})]$ cover $t^{\prime}$, so $\mathcal{I}$ cannot be active at $t^{\prime}$.
Initialization.
At the start, $\tilde{\mathbf{s}}=[(\mathbf{x}_{0},0)]$ and $\mathbb{T}$ is the given constraint set; by assumption it is feasible, so (I1) holds. No progress has been assigned, so (I2) and (I3) are vacuous.
Preservation under one extension.
Suppose the algorithm chooses a remaining reachability $\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)$ at $(\mathbf{x},t)$ and calls Algorithm 2. By construction, the routine computes the largest currently feasible time window $t_{\min},t_{\max}]$ consistent with $\mathbb{T}$, and samples a candidate $\mathbf{x}^{\prime}\models\mu$ together with a predicted arrival time $t^{\prime}$. If $t^{\prime}>t_{\max}$ or the residual feasible window is covered by conflicts, the attempt is discarded; otherwise the routine returns the earliest nonconflicting time $t_{\mathrm{new}}\in[\max\{t^{\prime},t_{\min}\},t_{\max}]\setminus\mathcal{O}$. Therefore, with $t^{*}:=t_{\mathrm{new}}$, the augmented constraints $a_{\Lambda}\leq t^{*},\ b_{\Lambda}\geq t^{*}$ are consistent with the current $\mathbb{T}$, establishing (I1) and (I2) at the child node. Moreover, for every invariance with determined start time that is violated by $\mathbf{x}^{\prime}$, UpdateConstraint inserts $d_{\Lambda}<t^{*}$; since Algorithm 2 avoided $\mathcal{O}$ and $t^{*}$ lies strictly to the right of the current lower bound for $d_{\Lambda}$, the new inequality is compatible with $\mathbb{T}$. Hence (I3) holds as well.
Termination case and witness construction.
When the algorithm returns, $\mathbb{P}^{\mathcal{R}}_{\mathrm{rem}}=\varnothing$ and the final constraints are $\mathbb{T}_{f}$ with feasible set $\mathcal{F}_{f}\neq\varnothing$ by (I1). Pick any $\boldsymbol{\lambda}\in\mathcal{F}_{f}$. By (I2), for every assigned reachability with waypoint time $t_{i}$, we have $t_{i}\in[a_{\Lambda}(\boldsymbol{\lambda}),b_{\Lambda}(\boldsymbol{\lambda})]$, hence $\mathbf{s}_{0}\models\mathcal{R}(a_{\Lambda}(\boldsymbol{\lambda}),b_{\Lambda}(\boldsymbol{\lambda}),\mu)$. For any invariance $\mathcal{I}(a_{\Lambda},b_{\Lambda},\mu^{\prime})$ and any waypoint time $t_{i}$ that lies in its active window $a_{\Lambda}(\boldsymbol{\lambda}),b_{\Lambda}(\boldsymbol{\lambda})]$, there are two possibilities: (i) $\tilde{\mathbf{x}}_{i}\models\mu^{\prime}$, in which case the waypoint trivially satisfies the invariance; or (ii) $\tilde{\mathbf{x}}_{i}\nvDash\mu^{\prime}$, in which case the start of this invariance has already been determined by its preceding reachability, and (I3) guarantees that $t_{i}$ cannot lie in the interval (because $d_{\Lambda}<t_{i}$ was added). Either way, the stated waypoint-level invariance condition holds. This proves the lemma. ∎
Remark 3 (Scope of the guarantee).
The lemma certifies that, under some feasible time-variable assignment, all reachability progresses are satisfied and no invariance progress is violated at the returned waypoint times. It does not claim invariance satisfaction at every intermediate time step between waypoints.
G.3 Proof of Theorem 1
Proof.
Let $(\mathbb{P}_{\varphi},\mathbb{T}_{\varphi})$ be the progress/constraint pair returned by the Semantics-based Decomposition, and let $\tilde{\mathbf{s}}=(\tilde{\mathbf{x}}_{0},t_{0})\,(\tilde{\mathbf{x}}_{1},t_{1})\,\dots\,(\tilde{\mathbf{x}}_{n},t_{n})$ with $0=t_{0}\leq\cdots\leq t_{n}$ and final constraint set $\mathbb{T}_{f}$ be the output of the Progress Allocation algorithm. By assumption, $\mathcal{F}_{f}\neq\varnothing$. Fix any feasible assignment $\boldsymbol{\lambda}^{\star}\in\mathcal{F}_{f}$ and let $\mathbf{s}_{0}=\mathbf{x}_{0}\mathbf{x}_{1}\dots\mathbf{x}_{T}$ ($T\geq t_{n}$) be the signal induced by the trajectory $\boldsymbol{\tau}$ returned by the Trajectory Generation module, which (by assumption) satisfies: (i) $\mathbf{x}_{t_{i}}=\tilde{\mathbf{x}}_{i}$ for all $i$, and (ii) for every invariance $\mathcal{I}(a_{\Lambda},b_{\Lambda},\mu)$, $\forall t\in[a_{\Lambda}(\boldsymbol{\lambda}^{\star}),\,b_{\Lambda}(\boldsymbol{\lambda}^{\star})]$ we have $\mathbf{x}_{t}\models\mu$.
Reachability progresses.
By construction of the allocation phase and by Lemma 2, for every $\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)\in\mathbb{P}_{\varphi}$ there exists an index $i$ such that (a) $\tilde{\mathbf{x}}_{i}\models\mu$ and (b) $t_{i}\in[a_{\Lambda}(\boldsymbol{\lambda}^{\star}),\,b_{\Lambda}(\boldsymbol{\lambda}^{\star})]$, because $\mathbb{T}_{f}$ contains the constraints $a_{\Lambda}\leq t_{i}$ and $b_{\Lambda}\geq t_{i}$, which are satisfied by $\boldsymbol{\lambda}^{\star}$. Since the generated signal visits all waypoints, $\mathbf{x}_{t_{i}}=\tilde{\mathbf{x}}_{i}$, hence $\mathbf{s}_{0}\models\mathcal{R}(a_{\Lambda}(\boldsymbol{\lambda}^{\star}),b_{\Lambda}(\boldsymbol{\lambda}^{\star}),\mu)$.
Invariance progresses.
By property (ii) of the generated trajectory, for every $\mathcal{I}(a_{\Lambda},b_{\Lambda},\mu)\in\mathbb{P}_{\varphi}$ we have $\forall t\in[a_{\Lambda}(\boldsymbol{\lambda}^{\star}),\,b_{\Lambda}(\boldsymbol{\lambda}^{\star})]$: $\mathbf{x}_{t}\models\mu$, i.e., $\mathbf{s}_{0}\models\mathcal{I}(a_{\Lambda}(\boldsymbol{\lambda}^{\star}),b_{\Lambda}(\boldsymbol{\lambda}^{\star}),\mu)$.
Conclusion via decomposition soundness.
We have shown that for the same feasible assignment $\boldsymbol{\lambda}^{\star}$,
$\forall\mathcal{P}\in\mathbb{P}_{\varphi}:\quad\mathbf{s}_{0}\models\mathcal{P}(\boldsymbol{\lambda}^{\star}).$
Therefore, by Lemma 1, $\mathbf{s}_{0}\vDash\varphi$. This proves the theorem. ∎